OpenStudy (anonymous):

if 27% of the group voted to change its bylaws and 43% voted against the change, how many people did not vote? the group has 810 members

ganeshie8 (ganeshie8):

\(\textbf{percent of people voted = 27% + 43% = 70%} \)

ganeshie8 (ganeshie8):

so... . what percent of people didnt vote ?

OpenStudy (anonymous):

(.27)+(.43)= .7 1-(.7)= .3 \[810\times.3=243\]

OpenStudy (anonymous):

is this the right answer ashley?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks! im taking a test right now thats why:)